An AC voltage of the form Δv = 100.0 sin (100.0t), where Δv is in volts and t in seconds), is applied to a series RC circuit. The value of the resistor is 1.00kΩ and the value of the capacitor is 10.0 μF. Determine the impedance for this circuit and then use it to determine the rms current flowing in this circuit.
Answer #1

Impedence Z^2 = R^2 + XC^2

where Xc = 1/wC

where W = 2pif here = W = 100 rad/s

R is resistance and C is capacitance

so

Xc = 1/(100 * 10 e-6) = 1000ohms

R = 1000 ohms

so

impedence Z^2 = 1000^2 + 1000^2

Z = 1.414 * 1000

Z = 1414 ohms-----------<<<<<<<<<Answer to Impedence

-----------------------------------------

Current I = V/Z

I = 100 /(1414)

I = 0.0707 Amps